First order reaction the rate in 0.1 s. Thus the 20g reactant when reduce to 5g calculate time?

Grade:11

1 Answers

Ankit
6 years ago
Given that Initial quantity, [R]o= 20gFinal quantity, [R] = 5g Rate constant, k = 1 x 10−1 s−1Formula of 1st order reaction,. . . . . .K=(2.303/t)(log[R]o/log[R]). . . . . . . . . . . . . . . . . . . Plug the values we get Value of log 20/5 = log 20 – log 5 = 1.3010-0.6990. . . . . . . .= (0.6020). . . . . . . . . . After applying formula k= (2.303/1*10-1)(0.6020). . . . . . . . . .=(23.03*0.6020) After calculation we get 13.864 sec . . . . . . .Hence, it will take 13.864 sec to reduce 20 g of this reactant to 5 g. . . . . . . . Thank u